### v2

@@ 1 Promising Age of Quantum Computing
1.3 Quantum Computing in a nutshell: A binary bit that may take on values ‘0’ or ‘1’ is the basic unit of information of conventional computing. Quantum Computation and Information uses qubits as fundamental unit of information and, differently from classical bits, they can not only acquire either value ‘0’ or ‘1’, but even ‘0’ and ‘1’ at the same time. A simple mathematical representation of a qubit, in the computational basis ${|0\rangle,|0\rangle}$, is conventionally given as:
@@
where $a$ and $b$ are complex amplitudes, $a,b\in\mathcal{C}$, superimposing the states ‘0’ and ‘1’ [2]. Using qubit superposition, a vast computational space becomes available, allowing for the solution of several problems of extreme complexity. Even a very limited number of qubits, $N$ can be used to solve problems that are intractable with classical computers, thanks to the rapidly expanding computational domain as an exponential function ($2^{N}$) of the total number of qubits.
@@ 4 Modern Cryptography: From Quantum to Post-Quantum
4.1 Quantum Key Distribution: Since quantum algorithms (e.g., Shor’s) endanger the classical cryptographic key exchange techniques (e.g., Diffie–Hellman key exchange), the Quantum Key Distribution (QKD) has been developed as a potential remedy to resist quantum attacks. The distribution of quantum keys, or QKD, is a powerful method for securing data using quantum computing devices  [4]. QKDs are primarily based on hybrid quantum mechanisms and are virtually impenetrable. An essential tenet of quantum encryption is the no-cloning hypothesis [47], which asserts that a flawless replica of a quantum network or its configurations cannot be created. Any effort to gain insight into quantum communication would result in the introduction of noise that can be readily identified by either the key distributor or the recipient. QKD techniques can be categorized by their approach to aspects such as application modification, encoding/decoding, and quantum channel implementations during transmissions.

### v3

@@ 1 Promising Age of Quantum Computing
1.3 Quantum Computing in a nutshell: A binary bit that may take on values ‘0’ or ‘1’ is the basic unit of information of conventional computing. Quantum Computation and Information uses qubits as fundamental unit of information and, differently from classical bits, they can not only acquire either value ‘0’ or ‘1’, but even ‘0’ and ‘1’ at the same time. A simple mathematical representation of a qubit, in the computational basis $\{|0\rangle,|1\rangle\}$, is conventionally given as:
@@
where $a$ and $b$ are complex amplitudes ($a,b\in\mathcal{C}$) superimposing the states ‘0’ and ‘1’ [2], and preserving probability interpretation of quantum state, i.e., they need to verify the condition $|a|^{2}+|b|^{2}=1$. The symbol |$\rangle$| ⟩ (ket) indicates that the bit of information is encoded in a quantum state, exploiting one of its physical degree of freedom. Using quantum superposition, a vast computational space becomes available allowing to solve problems of extreme complexity [INCLUDE A REFERENCE HERE on quanutm computing basics]. Even a very limited number of qubits, $N$ can be used to solve problems that are intractable with classical computers, thanks to the rapidly expanding computational domain as an exponential function ($2^{N}$) of the total number of qubits.
@@ 4 Modern Cryptography: From Quantum to Post-Quantum
4.1 Quantum Key Distribution: Since quantum algorithms (e.g., Shor’s) endanger the classical cryptographic key exchange techniques (e.g., Diffie–Hellman key exchange), the Quantum Key Distribution (QKD) has been developed as a potential remedy to resist quantum attacks. QKDs protocols are always hybrid, i.e., they rely on both quantum and classical communications to implement a virtually impenetrable crypto-system that promises to protect the privacy of communication even against attacks conducted by quantum computers. An essential tenet of such an approach to encryption is roothed in quantum physics and takes the shape of the no-cloning theorem [47], which asserts that a flawless replica of arbitrary (i.e. non-orthogonal) quantum states cannot be created without corrupting the probed quantum states. That means that any effort to gain insight into such non-orthogonal quantum states, used to encode information, would results in the introduction of noise that can be readily identified by either the key distributor or the recipient (the parties, conventionally Alice and Bob). Such a mechanism allows the parties to quantify the amount of information potentially eavesdropped during the quantum communication, and then apply classical protocols of error correction and privacy amplification that can reduce the eavesdropper’s knowledge on the shared key to a negligible amount. This second part of the cryptosystem is usually called classical communication phase phase.